OpenStudy (anonymous):

need the solutions please. √(y+10)=y-2

OpenStudy (anonymous):

\[\sqrt{y+10} -y = -2\]

OpenStudy (anonymous):

i was getting there ")

OpenStudy (anonymous):

ok maybe not

OpenStudy (callisto):

\[\sqrt{y+10}-y = -2\]\[ y+10 = (y-2)^{2}\]\[ y^{2}-5y-6 = 0\]y=6 or y=-1

OpenStudy (anonymous):

sorry, my mistake :(

OpenStudy (anonymous):

its alright